OpenStudy (anonymous):

Consider the following functions f(x)= (7x+8)/(x+3) and g(x)= (3x-8)/(7-x) (a) Find g(g(x)) (b) Find g(f(x)) (c) Determine whether the functions f and g are inverses of each other.

OpenStudy (anonymous):

bunch o algebra for this one

OpenStudy (anonymous):

is the first one \(g(g(x))\) or is it \(f(g(x))\) ?

OpenStudy (anonymous):

f(g(x)) sorry

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok ready?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f(g(x))=f\left(\frac{3x-8}{7-x}\right)\] is the first step

OpenStudy (anonymous):

then where you see an \(x\) in \(f(x)=\frac{7x+8}{x+3}\) replace it by \(\frac{3x-8}{7-x}\) to get \[f(g(x))=\frac{7\left(\frac{3x-8}{7-x}\right)+8}{\frac{3x-8}{7-x}+3}\]

OpenStudy (anonymous):

then simplify the compound fraction by multiplying top and bottom by \(7-x\)

OpenStudy (anonymous):

the first step is to write \[\frac{7(3x-8)+8(7-x)}{3x-8+3(7-x)}\]then multiply out, cancel etc

OpenStudy (anonymous):

Oh ok gotcha!

OpenStudy (anonymous):

if you do it carefully, you will see that there is a whole mess of cancellation you will just get \(x\) which shows that \[f(g(x))=x\] which is a good indication that \(f\) and \(g\) are inverse functions